# Distributional Regression with Tabular Foundation Models: Evaluating Probabilistic Predictions via Proper Scoring Rules

### 09 Mar 2026 v1

@@ 3 Results
We adapt the finetuning code from github to start from the realTabPFNv2.5 checkpoint and implement custom loss functions (Beta energy score (with CRPS as special case), CRLS, and others).
We fit models on a diverse set of openML datasets for regression as shown in the various result tables. For each dataset we limit the amount of samples to randomly selected 3000 samples (random subsampling). We then perform 5 fold cross-validation on each subsampled dataset so that we can report performance on holdout data. On each split we evaluate a baseline model (always the realTabPFNv2.5) against another model (either a finetuned model or TabICLv2), so that we get paired performance metrics (per split). We report several performance metrics ($MAE$, RMSERMSE, $R^{2}$, $CRPS$, CRLSCRLS, $IntervalScore@95$). All metrics except R2 are better when they are lower. The improvement in $R^{2}$ is reported as absolute difference in $\Delta R^{2}*100$pp (to avoid statistical problems with reporting percent changes on fractions). The other improvements are reported as relative change from the baseline.
For simple interpretation we adopt the sign convention that a positive change is an improvement in the respective metric. For the metrics we first compute the change in the metric (at same split) and then compute the average afterwards.
Table 1: Finetuned TabPFN Regressor Hyperparameters
@@
3.1 Beta(1.8)-Energy Score Finetuning realTabPFNv2.5
@@
We optimize with $\beta=1.8$ energy score and apply early stopping on MAE (changing the metric for early stopping showed strongly varying performance). We find improved performance in terms of $R^{2}$:
Table 2: Comparative Performance and Uncertainty Metrics (Finetuned vs. Base)
Metric Improvement (mean $\pm$ std) Median W/L/T
MAE +4.28% $\pm$ 10.06% +1.46% 61/21/8
RMSE +2.16% $\pm$ 4.39% +0.80% 61/24/5
$R^{2}$ +1.47pp $\pm$ 3.73pp +0.25pp 52/19/19
CRPS +2.76% $\pm$ 6.38% +0.79% 61/23/6
IntervalScore95 +2.24% $\pm$ 13.86% +0.94% 55/33/2
Table 3: Improvement by Dataset: Finetuned vs. Base
Dataset MAE Imp. RMSE Imp. $R^{2}$ Imp. CRPS Imp. IS95 Imp.
Abalone -0.03% +0.47% +0.39pp +0.78% +0.60%
Ailerons -0.04% -0.24% -0.06pp +0.18% -0.30%
Bike Sharing +1.49% +1.11% +0.13pp +1.08% -0.49%
Black Friday +0.16% +0.30% +0.30pp +0.25% -0.74%
Brazilian Houses -7.33% +3.43% +6.28pp +0.99% +5.64%
California Housing +0.91% +1.10% +0.48pp +1.25% +1.37%
Cpu Small +1.38% -0.12% -0.00pp -0.03% +0.92%
Diamonds +0.61% +1.26% +0.05pp +0.53% -1.46%
Digits +34.54% +6.29% +0.01pp +8.27% +13.74%
Elevators +0.02% +0.20% +0.03pp +0.21% +0.82%
House 16H -0.39% -0.06% -0.08pp -0.47% -0.69%
Kin8nm +1.83% +2.05% +0.30pp +2.01% +5.18%
Mercedes Benz +15.71% +6.94% +7.33pp +13.58% +7.27%
NYC Taxi +8.64% +4.73% +6.68pp +8.18% +3.50%
Protein +1.44% +0.26% +0.25pp +0.51% +1.12%
Pol +8.31% +3.30% +3.71pp +5.75% +9.80%
Puma8NH +7.87% +7.95% +0.77pp +6.85% +7.06%
Sulfur +1.85% -0.04% -0.02pp -0.17% -13.03%
Figure 4: Impact of beta=1.8 energy score finetuning on realTabPFNv2.5
3.2 CRLS-Finetuning realTabPFNv2.5
@@
We implement the CRLS as a loss in tabpfn and use it here also as evaluation metric for early stopping.
Table 4: Overall Performance and Uncertainty Improvement (Finetuned vs. Base)
Metric Improvement (mean $\pm$ std) Median W/L/T
MAE +0.65% $\pm$ 10.13% +0.21% 54/41/5
RMSE +1.67% $\pm$ 4.60% +0.96% 73/25/2
$R^{2}$ +0.94pp $\pm$ 2.40pp +0.33pp 64/20/16
CRLS +2.27% $\pm$ 7.00% +1.47% 79/18/3
IntervalScore95 +3.87% $\pm$ 13.15% +1.76% 75/22/3
Table 5: Detailed Dataset Improvements: Finetuned vs. Base
Dataset MAE Imp. RMSE Imp. $R^{2}$ Imp. CRLS Imp. IS95 Imp.
Abalone -0.76% +0.42% +0.35pp +1.52% +2.58%
Ailerons +1.01% +0.65% +0.17pp +1.52% +0.86%
Ames Housing -0.21% +1.41% -0.23pp -0.75% -7.59%
Bike Sharing +1.86% +1.88% +0.21pp +1.19% +1.59%
Black Friday +0.23% +0.61% +0.61pp +1.09% +0.67%
Brazilian Houses -24.19% +3.67% +6.72pp -2.45% +23.84%
California Housing +0.92% +1.94% +0.74pp +2.48% +3.10%
Cpu Small +0.69% +0.05% +0.00pp +1.14% +1.45%
Diamonds +0.60% +1.27% +0.05pp +1.01% +1.15%
Digits +28.23% +7.48% -0.02pp +10.91% +3.09%
Elevators -0.54% +0.13% +0.02pp +0.34% +0.67%
House 16H -1.88% +0.42% +0.27pp +0.20% +0.61%
House Sales -0.71% -1.35% -0.34pp +0.22% -1.02%
Kin8nm -0.39% -0.77% -0.12pp -0.32% +0.97%
Mercedes Benz +3.46% +1.14% +1.22pp +2.12% +1.31%
NYC Taxi +4.71% +3.04% +3.90pp +4.19% +2.08%
Protein -1.47% +0.51% +0.45pp +1.43% +3.34%
Pol +7.37% +3.23% +3.64pp +8.15% +13.82%
Puma8NH +7.24% +7.41% +0.71pp +6.26% +6.37%
Sulfur -13.24% +0.36% +0.48pp +5.14% +18.50%
Figure 5: Impact of CRLS finetuning realTabPFNv2.5
3.3 CRLS performances of TabICLv2 versus realTabPFNv2.5
Table 6: Aggregate Performance and Uncertainty Improvement (TabICLv2 vs. realTabPFNv2.5)
Metric Improvement (mean $\pm$ std) Median W/L/T
MAE +1.41% $\pm$ 12.65% +0.16% 56/48/1
RMSE +1.85% $\pm$ 10.02% +0.96% 65/34/6
$R^{2}$ +3.22pp $\pm$ 17.03pp +0.17pp 56/27/22
CRLS +6.01% $\pm$ 10.95% +3.49% 84/19/2
IntervalScore95 +5.14% $\pm$ 14.18% +2.15% 70/35/0
Table 7: Dataset-Specific Improvement Metrics
Dataset MAE Imp. RMSE Imp. $R^{2}$ Imp. CRLS Imp. IS95 Imp.
Abalone +0.22% +0.92% +0.76pp +3.37% +4.94%
Ailerons -1.69% -1.01% -0.29pp +1.55% +0.23%
Ames Housing +0.37% +8.13% +32.38pp +4.95% -13.51%
Bike Sharing +0.21% -0.62% -0.06pp +2.83% -0.13%
Black Friday +0.59% +1.17% +1.18pp +0.34% -2.20%
Brazilian Houses -13.82% +3.41% +6.13pp +25.02% +31.86%
California Housing +4.87% +5.25% +1.91pp +8.33% +6.34%
Cpu Small +1.49% +0.08% +0.00pp +6.17% +3.52%
Diamonds +1.15% +1.62% +0.06pp +4.33% +4.93%
Digits +42.92% +21.15% +0.89pp +34.26% +43.52%
Elevators -0.20% +0.57% +0.10pp +1.55% -0.55%
House 16H +0.76% +2.92% +1.92pp +5.30% +5.42%
House Sales -5.00% -10.86% -2.37pp +0.11% -3.01%
Kin8nm +0.59% -0.29% -0.04pp +1.52% +3.26%
Mercedes Benz +18.83% +9.57% +9.92pp +20.31% +13.27%
NYC Taxi -7.26% +3.99% +9.25pp -5.47% -6.48%
Protein +2.63% +1.71% +1.55pp +4.28% +3.15%
Pol +10.39% +4.05% +4.56pp +11.22% +7.74%
Puma8NH -13.29% -13.86% -1.50pp -13.66% -10.89%
Sulfur -10.13% +1.27% +1.61pp +6.38% +16.39%
Visualizing Soil -4.08% -0.36% -0.32pp +3.50% +0.08%
Figure 6: Impact of changing from RealTabPFNv2.5 to TabICLv2: Even stronger outperformance of TabICLv2 over realTabPFNv2.5 on CRLS than seen before by finetuning realTabPFNv2.5 on CRLS
3.4 CRPS performances of realTabPFNv2.5 and TabICLv2
Table 8: Overall Model Improvement (TabICLv2 vs. realTabPFNv2.5)
Metric Improvement (mean $\pm$ std) Median W/L/T
MAE +1.41% $\pm$ 12.65% +0.16% 56/48/1
RMSE +1.85% $\pm$ 10.02% +0.96% 65/34/6
$R^{2}$ +3.22pp $\pm$ 17.03pp +0.17pp 56/27/22
CRPS +2.01% $\pm$ 10.17% +0.73% 59/45/1
IntervalScore95 +5.14% $\pm$ 14.18% +2.15% 70/35/0
Table 9: Dataset-Specific Performance Gains
Dataset MAE Imp. RMSE Imp. $R^{2}$ Imp. CRPS Imp. IS95 Imp.
Abalone +0.22% +0.92% +0.76pp +0.86% +4.94%
Ailerons -1.69% -1.01% -0.29pp -2.20% +0.23%
Ames Housing +0.37% +8.13% +32.38pp -1.81% -13.51%
Bike Sharing +0.21% -0.62% -0.06pp -0.75% -0.13%
Black Friday +0.59% +1.17% +1.18pp +0.61% -2.20%
Brazilian Houses -13.82% +3.41% +6.13pp +3.63% +31.86%
California Housing +4.87% +5.25% +1.91pp +5.84% +6.34%
Cpu Small +1.49% +0.08% +0.00pp +1.82% +3.52%
Diamonds +1.15% +1.62% +0.06pp +1.45% +4.93%
Digits +42.92% +21.15% +0.89pp +28.19% +43.52%
Elevators -0.20% +0.57% +0.10pp -0.17% -0.55%
House 16H +0.76% +2.92% +1.92pp +1.82% +5.42%
House Sales -5.00% -10.86% -2.37pp -4.79% -3.01%
Kin8nm +0.59% -0.29% -0.04pp +0.20% +3.26%
Mercedes Benz +18.83% +9.57% +9.92pp +20.91% +13.27%
NYC Taxi -7.26% +3.99% +9.25pp -8.96% -6.48%
Protein +2.63% +1.71% +1.55pp +2.79% +3.15%
Pol +10.39% +4.05% +4.56pp +6.73% +7.74%
Puma8NH -13.29% -13.86% -1.50pp -15.15% -10.89%
Sulfur -10.13% +1.27% +1.61pp +1.78% +16.39%
Visualizing Soil -4.08% -0.36% -0.32pp -0.56% +0.08%
4 Choice of Scoring Rules for Decision Making
The choice of a scoring rule matters because different scoring rules have different inductive biases Waghmare and Ziegel (2025); Merkle and Steyvers (2013). This is the reason why the scoring rule used for solving a business problem should be chosen carefully:
  • Oesterheld and Conitzer (2020) describes "Decision Scoring Rules"
  • Winkler et al. describe tailored scoring rules for decision making from a utility perspectiveJohnstone et al. (2011).
  • The very insightful introduction example in Johnstone et al. (2011), page 2, is about making "useful" probabilistic weather forecasts for a specific use-case (an orchard) and we drew a lot of inspiration reading it - this example is very intuitive we suggest to read itJohnstone et al. (2011).
  • Merkle et al.Merkle and Steyvers (2013) in their paper "Choosing a strictly proper scoring rule" describe an important effect: The "optimal forecaster" depends on the choice of a scoring rule (while all scoring rules are individually minimized by a forecast which is equal to the true distribution).
In other words, different scoring rules penalize different types of errors differently and thus the optimal forecast (the one which minimizes the expected score) can differ based on the scoring rule used. Merkle already points this out and finds that different forecasters are ranked very differently by different strictly proper scoring rules. They for example find in table 1 that the spearman correlation for the ranking of forecasters between the brier score and the log score is only 0.15. They then go on and investigate a parametrizable family of scoring rules and see how the ranking generally depends on the choice of the parameters.
For this article we follow this idea and port it to proper scoring rules which can be used in the regression setting. We choose the $\beta$-energy score family and show that the optimal model depends on the choice of $\beta$. To show this and for keeping the arguments as simple as possible we consider
4.1 Invesigating the effect of the choice of scoring rule on the optimal forecast
$\hat{p}_{i}(Y=y|X=x)=\delta(y-f_{i}(x)),$ (7)
$p_{g}(Y=y|X=x)=\delta(y-g(x)).$ (8)
We choose $x\in[-3,3]$ (np.linspace(-3,3,1000)).
  • $g(x)=\sin(x)$
  • Model A: Small consistent bias, $f_{A}(x)=g(x)+0.1$
  • Model C: Scaled version of the true function, $f_{C}(x)=0.8\cdot g(x)$
  • Model D: Shifted version of the true function, $f_{D}(x)=\sin(x+0.2)$
  • Model E: Simple baseline at zero, $f_{E}(x)=0$
A typical user would now wonder which model is best? To get an intuition, we look at the models in Figure 8.
Obviously the different models make different erros which can be seen in the absolute error plot in Figure 9. Some models make small errors everywhere (model A), some models make large errors in a small region (model B), some models make medium sized errors everywhere (model C) and some models make small errors everywhere but are shifted (model D).
Figure 9: Absolute errors for the toy models
Now we can apply different scoring rules (from the $\beta$-energy family) to see which model is optimal for which scoring rule. For a fixed choice of $\beta$ (one point on the x-axis in Figrue 10) the better model has the lower beta energy score.
Figure 10: Decision based on scoring rules for the toy models
We observe that while the model B is the best ranked model for each $\beta\lesssim 1$ for $1\lesssim\beta$ this changes and also other models switch rank at other values of $\beta$, see figure 11.
Figure 11: Ranking of toy models based on scoring rules
Fundamentally this behavior arises from different scoring rules rating the severity of errors in Figure 9 differently (and there is even scoring rules wich penalize errors depending on the direction of the error, see for example asymmetric CRPSIacopini et al. (2023) which may be important e.g. in finance where errors have different severity depending on the direction of the error: making 100 percent more money is less severe than loosing 100 percent of money).
For simple probabilistic point masses like shown here we can even make an analytical statement: The beta energy score with $\beta=1$ favors minimization of the mean absolute error (i.e. prefers estimating a median estimate) while for $\beta=2$ the scoring rule prefers a minimization of mean squared error.
Derivation.
where $X$ and $X^{\prime}$ are independent draws from $P$.
$\mathrm{ES}_{\beta}(\delta_{\hat{m}},\,y)=|\hat{m}-y|^{\beta}.$ (10)
The choice of the scoring rule determines how different types of errors are penalized and thus the optimal forecast (the one which minimizes the expected score) can differ based on the scoring rule used. This is true, despite every scoring rule being strictly proper (i.e. every scoring rule is minimized by the true distribution). The reason is that different scoring rules penalize different types of errors differently and thus the optimal forecast (the one which minimizes the expected score) can differ based on the scoring rule used.
We leave it as an open question how to best choose a scoring rule for general purpose tabular foundation model pretraining and how to enable adaptation to arbitrary business case with high severity for certain errors (like in finance or pharmaceutical production). We describe some possible routes to explore in the outlook section 7.
5 Discussion
Across the datasets evaluated, finetuning with beta-energy score ($\beta=1.8$) and CRLS both tend to improve realTabPFNv2.5 on the majority of metrics and datasets (as indicated by the W/L/T counts in Tables 2 and 4). However, improvements are modest in median terms (typically $2\%$) and vary considerably across datasets.
The aggregate results further suggest that TabICLv2 may be a stronger probabilistic model than realTabPFNv2.5 in terms of CRPS, CRLS, and Interval Score, with TabICLv2 winning on the majority of evaluated datasets (Tables 6, 8). The margin varies substantially across datasets (e.g. large gains on Mercedes Benz and Pol, but losses on Puma8NH and House Sales).
Additional experiments examining how finetuning performance scales with dataset size would be valuable to better understand whether the observed differences are robust across data regimes.
We illustrate in section 4.1 that the choice of scoring rule can change the ranking of forecasting models — at least in the considered toy setting of Dirac-delta forecasts over a smooth ground truth function. Under this simplification, the result follows analytically from the $\beta$-energy score reducing to an $L^{\beta}$ norm. Whether this ranking sensitivity carries over to actual trained distributional models in realistic tabular settings remains an open empirical question. That said, the theoretical argument and the empirical literature Waghmare and Ziegel (2025); Merkle and Steyvers (2013) both support the view that the choice of scoring rule for pretraining or finetuning a tabular foundation model is non-trivial and may affect which use-cases the model is best suited for Oesterheld and Conitzer (2020); Johnstone et al. (2011).
6 Limitations: Rare or Extreme Events
We want to highlight that optimal scoring rule estimation (minimizing/maximizing the expected value of a scoring rule) is inherently not well suited for eliciting properties which are only realized in rare events (fall in the area where no data was observed or in 1D are in the tail of a distribution). This was well argued in recent years for the case of one dimensional distributions but equally carries over to higher dimensions. Lerch et al. (2017); Brehmer and Strokorb (2019) Inherently, this problem is related to epistemic uncertainty (uncertainty due to not having training data). Based on this insight also PFNs (trained to optimize mean scoring rules) should be applied with care if risk associated mainly with rare (tail) events or events outside the support of the examples in the supplied context.

### 13 Mar 2026 v2

@@ 5 Results
We adapt the TabPFN fine-tuning code from GitHub, starting from the realTabPFNv2.5 checkpoint, and implement custom loss functions (the $\beta$-energy score, with CRPS as the $\beta=1$ special case, and CRLS).
We fit models on a diverse set of OpenML datasets for regression:
  • Abalone
  • Ailerons
  • Ames Housing
  • Bike Sharing
  • Black Friday
  • Brazilian Houses
  • California Housing
  • Cpu Small
  • Diamonds
  • Digits
  • Elevators
  • House 16H
  • House Sales
  • Kin8nm
  • Mercedes Benz
  • NYC Taxi
  • Protein
  • Pol
  • Puma8NH
  • Sulfur
Each dataset is randomly subsampled to at most 3000 instances. We then apply 5-fold cross-validation on each subsampled dataset, yielding holdout performance estimates on all available data. On each fold, the baseline model (realTabPFNv2.5) and the comparison model (either a fine-tuned variant or TabICLv2) are evaluated on the same held-out split, producing paired performance metrics.
We report six performance metrics: $\mathrm{MAE}$, $\mathrm{RMSE}$, $R^{2}$, $\mathrm{CRPS}$, $\mathrm{CRLS}$, and $\mathrm{IS}_{95}$ (Interval Score at the 95% level). All metrics are lower-is-better, with the exception of $R^{2}$. Improvements in $R^{2}$ are expressed as the absolute difference $\Delta R^{2}\times 100$ percentage points (pp), avoiding the statistical ambiguity of relative changes on a fraction. Improvements in all other metrics are reported as relative changes with respect to the baseline. Throughout, we adopt the sign convention that a positive value denotes an improvement over the baseline, regardless of the metric’s direction. For aggregated results, per-fold improvements are computed first and then averaged across folds and datasets.
The hyperparameters used for fine-tuning the TabPFN regressor are shown in Table 2.
Table 2: Fine-Tuned TabPFN Regressor Hyperparameters. The built-in loss weights (crps_loss_weight etc.) are explicitly set to 0.0 to override the model’s default loss configuration; all weight is instead assigned to the custom scoring-rule implementation.
@@
We fine-tune TabPFN with multiple scoring rules; here we report results for CRLS and the $\beta=1.8$ energy score, evaluating each variant against the base model (realTabPFNv2.5).
@@
  • We optimize with $\beta=1.8$ energy score and apply early stopping on MAE (changing the metric for early stopping showed strongly varying performance, because the early stopping metric is also involved in model selection through picking specific checkpoints).
@@
  • We implement CRLS as a loss function in TabPFN and use it also as the early-stopping metric.
  • We additionally report TabICLv2 results against the base model to contextualize the fine-tuned variants relative to another strong baseline.
Aggregated paired performance improvements above the baseline are reported in Table 3.
Table 3: Aggregated paired performance improvement across all model comparisons. All improvements are relative to the realTabPFNv2.5 baseline. Dashes indicate metrics not evaluated for that comparison. Bold values indicate the best performer for that specific column/metric combination. W/L/T counts the number of datasets where the model exceeded the baseline by more than a small $\varepsilon$ (Win), the baseline was better by more than $\varepsilon$ (Loss), or the absolute difference was at most $\varepsilon$ (Tie).
Beta(1.8)-Energy FT CRLS FT TabICLv2
Metric mean $\pm$ std Median W/L/T mean $\pm$ std Median W/L/T mean $\pm$ std Median W/L/T
$\mathrm{MAE}$ +4.28% $\pm$ 10.06% +1.46% 61/21/8 +0.65% $\pm$ 10.13% +0.21% 54/41/5 +1.41% $\pm$ 12.65% +0.16% 56/48/1
$\mathrm{RMSE}$ +2.16% $\pm$ 4.39% +0.80% 61/24/5 +1.67% $\pm$ 4.60% +0.96% 73/25/2 +1.85% $\pm$ 10.02% +0.96% 65/34/6
$R^{2}$ +1.47pp $\pm$ 3.73pp +0.25pp 52/19/19 +0.94pp $\pm$ 2.40pp +0.33pp 64/20/16 +3.22pp $\pm$ 17.03pp +0.17pp 56/27/22
$\mathrm{CRPS}$ +2.76% $\pm$ 6.38% +0.79% 61/23/6 +2.01% $\pm$ 10.17% +0.73% 59/45/1
$\mathrm{CRLS}$ +2.27% $\pm$ 7.00% +1.47% 79/18/3 +6.01% $\pm$ 10.95% +3.49% 84/19/2
$\mathrm{IS}_{95}$ +2.24% $\pm$ 13.86% +0.94% 55/33/2 +3.87% $\pm$ 13.15% +1.76% 75/22/3 +5.14% $\pm$ 14.18% +2.15% 70/35/0
The results on the dataset level are reported in the appendix 8.2.
The empirical results demonstrate a clear performance trade-off between point-estimate accuracy and probabilistic calibration:
Point Estimation Superiority:
The Beta(1.8)-Energy FT variant is the strongest performer for absolute error reduction. It achieves the highest mean improvement in $\mathrm{MAE}$ ($4.28\%$) and $\mathrm{CRPS}$ ($2.76\%$), suggesting it is the most effective model when the primary objective is minimizing the distance between predictions and ground truth.
Consistency and Robustness:
CRLS FT exhibits high consistency across diverse datasets. While its mean improvements are more conservative, it secures the highest number of wins in $\mathrm{RMSE}$ ($73/25/2$) and $R^{2}$ ($64/20/16$). This indicates that the CRLS fine-tuning objective successfully mitigates the impact of outliers, offering a more stable "frequentist" profile.
Probabilistic Calibration:
TabICLv2 dominates the interval-based and probabilistic scoring metrics. It provides the most significant gains in $\mathrm{CRLS}$ ($6.01\%$) and $\mathrm{IS}_{95}$ ($5.14\%$). Furthermore, its superior mean $R^{2}$ improvement ($3.22\text{pp}$) suggests a high capacity for capturing complex, non-linear dependencies while maintaining well-calibrated uncertainty quantification.
We see that on average Beta(1.8)-Energy FT is beneficial for tasks requiring the most accurate point predictions in terms of MAE, while TabICLv2 is the superior choice for applications where high-quality confidence intervals and density estimation are critical or when a high mean $R^{2}$ is to be achieved. However, when median $R^{2}$ performance is to be optimized, fine-tuning with more tail-sensitive metrics like CRLS is beneficial.
Further differences between TabICLv2 and realTabPFNv2.5 may be due to other factors like
  • different prior data generating processes
  • different architecture with different memory scaling inducing different inductive biases
We note that frequent differences in $\mathrm{RMSE}$ and $R^{2}$ performance are due to relative improvement aggregation and different noise levels across datasets. Since $R^{2}$ is a scale-independent metric (effectively a per-dataset scaled $\mathrm{RMSE}$), it provides a more balanced view of performance across tasks with varying target variances, whereas mean improvements in $\mathrm{RMSE}$ can remain sensitive to the specific error scales of the underlying datasets.
5.1 TabICLv2 vs RealTabPFNv2.5 Single Dataset Results
The per-dataset results are shown in the appendix. The performance difference between TabICLv2 and realTabPFNv2.5 is heavily concentrated in specific datasets, revealing the strengths and edge cases of each PFN:
Point Estimation (MAE & RMSE):
The most substantial gains are observed in the Digits ($42.92\%$ MAE) and Mercedes Benz ($18.83\%$ MAE) datasets, where TabICLv2 appears to capture the underlying structure significantly better than the baseline. Conversely, TabICLv2 struggles with Puma8NH ($13.29\%$ MAE) and Brazilian Houses ($13.82\%$ MAE). Interestingly, in Brazilian Houses, the model loses in MAE but wins in RMSE ($3.41\%$), suggesting that while its median predictions may be further off, it is better at avoiding the large squared errors that penalize RMSE.
Goodness-of-Fit ($R^{2}$):
The improvement in $R^{2}$ is dominated by Ames Housing ($32.38\text{pp}$), which is a massive outlier compared to the rest of the benchmark. Significant positive shifts also occur in Mercedes Benz ($9.92\text{pp}$) and NYC Taxi ($9.25\text{pp}$). These high-impact gains are contrasted by minor losses in the House Sales ($2.37\text{pp}$) and Puma8NH ($1.50\text{pp}$) datasets, indicating that TabICLv2’s inductive bias is particularly potent for high-dimensional feature spaces like housing and automotive pricing.
Likelihood Calibration (CRLS):
CRLS improvements are nearly universal, but the magnitude is extreme in Digits ($34.26\%$) and Brazilian Houses ($25.02\%$). Even in datasets where point metrics are neutral or slightly negative, such as Cpu Small and Sulfur, the CRLS improves by $6.17\%$ and $6.38\%$ respectively. This demonstrates that even if the prediction centers are similar, TabICLv2 is far superior at modeling the shape of the uncertainty.
Global Error (CRPS):
The CRPS metric reinforces the dominance of TabICLv2 on Digits ($28.19\%$) and Mercedes Benz ($20.91\%$). Notably, in NYC Taxi, the model suffers a $8.96\%$ drop in CRPS, which is its second-largest loss after Puma8NH ($15.15\%$). This suggests that for certain temporal or high-frequency spatial datasets, the realTabPFNv2.5 prior remains highly competitive.
Interval Estimation (IS95):
The Brazilian Houses dataset shows the most dramatic improvement here ($31.86\%$), alongside Digits ($43.52\%$) and Sulfur ($16.39\%$). The only major degradation occurs in Ames Housing ($13.51\%$) and Puma8NH ($10.89\%$). The Ames result is particularly curious: while the model provides the best $R^{2}$ (explained variance) for this dataset, it provides poor interval coverage, suggesting the model is “overconfident” about its high-accuracy predictions on this specific task. This illustrates a general pitfall: a model can sharpen its point predictions (improving $R^{2}$) while simultaneously narrowing its predictive intervals too aggressively, producing intervals that are too tight and thus under-cover the true outcome. It is precisely this kind of trade-off that motivates reporting both point-estimate metrics and proper scoring rules such as the Interval Score side by side.
Summary: Overall, fine-tuning realTabPFNv2.5 as well as switching to TabICLv2 has a positive impact on several performance metrics compared to the baseline. However, as seen in Table 3, each model exhibits distinct strengths aligned with its training objective or coming from different pretraining datasets and architecture. Fine-tuning with beta-energy score ($\beta=1.8$) and CRLS both improve realTabPFNv2.5 on the majority of metrics and datasets (as indicated by the W/L/T counts in Tables 6 and 8), though improvements are modest in median terms (typically $2\%$) and vary considerably across datasets.
TabICLv2 appears to be a stronger probabilistic model than realTabPFNv2.5 in terms of CRPS, CRLS, and $\mathrm{IS}_{95}$, winning on the majority of evaluated datasets (Tables 10, 12). The margin varies substantially across datasets (e.g. large gains on Mercedes Benz and Pol, but losses on Puma8NH and House Sales). We attribute these complementary strength profiles in part to the fundamental inductive biases imposed through different scoring-rule-based loss functions: as discussed in section 4, optimizing a Beta-Energy score with $\beta$ closer to $2$ favors MSE minimization, and training with any one of the infinitely many possible scoring rules dictates how specific approximation errors are penalized.